Fermi National Laboratory

Accelerator Update for July 09, 2008 - July 11, 2008


Wednesday July 9

The day shift began with Operations monitoring the Recycler (RR) stash (159.7E10), with the Antiproton Source stacking (67.0E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) waiting for the next store, and with the SciBooNE / MiniBooNE (BNB) and Meson MTest experiments taking beam.

At 8:40 AM, the P1 transfer line beam permit tripped off due to low current on a power supply (I:VT701). Everything reset.

Operations put the Tevatron (TeV) through a wet squeeze at 8:55 AM.

Operations put the TeV into shot setup at 9:45 AM.

At 9:55 AM, Operations reported a problem with the TeV sector B3 BPMs. An expert investigated.

At 10:35 AM, the TeV suffered an abort at sectors A1, B1, and D0 during the final proton load.

Pbar experts reported at 11:01 AM, that power supply experts had completed their repair on Sweeper magnet “A.”

Recycler experts began a study around 11:20 AM.

Operations established store 6278 at 11:33 AM with an initial luminosity of 214.46E30.

At 12:33 PM, the water experts were at MI-60 switching to the 60 HP pumps, but left both of the pond pumps on too long and tripped off the LCW pumps and the MI Power Supply Interlock loop. This also caused the Recycler correctors to trip off, which dumped the stash (4.5E10).

Operations brought MI back on. Pbar resumed stacking at 1:06 PM.

At 2:29 PM, Operations turned off the MTest beam due to rising LCW temperatures at MI-10. The Run Coordinator said the experiment could take beam tonight for what they missed during the day.

Operators reported at 3:35 PM that they had search and secured NuMI (MI-65 and MINOS).

Operations reported at 4 PM that the Recycler had stashed antiprotons three times during the day shift.

Operations reported at 4:33 PM that they were preparing to send beam to NuMI for target scans.

At 4:47 PM, Operations reported that a NuMI Lambertson (LAM61) was tripping off on a ground fault indication. Operations called in an expert.

A power supply expert reported at 5:28 PM that LAM61 did not have a ground fault. He thought the problem was with the ground fault detector. He bypassed the alarm so that NuMI could complete the target scan.

Operations established beam to NuMI at 6:10 PM.

Operations established beam to MTest at 7:21 PM. Earlier they had disabled beam to BNB due to LCW temperature problems.

At 8:32 PM, the Linac beam intensity dropped by 50%. Operators investigated and discovered that the Linac buncher (IPA2) had tripped off. It reset.

Operations stopped sending beam to NuMI at 9 PM.

Operations resumed sending beam to BNB around 10:20 PM

Operations prepared MI-65 and MINOS for access at 11:51 PM.

Operations reported at 11:59 PM that the Recycler had stashed six times during the evening shift.

Thursday July 10

The midnight shift began with Operations monitoring store 6278 (60.20E30) and the stash (290.5E10), with Pbar stacking (24.9E10), and with BNB and MTest taking beam.

At 1:05 AM, three Pbar beam valves (BV728, BV 309, and BV210D) closed. The vacuum looked nominal. Everything reset, eventually.

At 3:16 AM, a Linac Klystron RF station (KRF1) tripped off for the first of two times in 15 minutes. It reset.

At 4:27 AM, the P1 transfer line tripped off due to an incorrect setting on I:VT701. Everything reset.

Operations terminated store 6278 at 5:32 AM.

Operations put the TeV into shot setup at 5:45 AM.

Operations disabled the MTest beam at 5:49 AM.

At 6:43 AM, Operations disabled BNB beam due to the high LCW temperature at MI-10.

At 6:46 AM, the TeV suffered a sector E4 quench.

Operations resumed sending beam to BNB at 7:02 AM.

Operations reported at 7:54 AM that the Recycler had stashed three times during the midnight shift.

Operations prepared the TeV sectors A-E for an access at 9:18 AM.

At 9:26 AM, Operations reported that the TeV sector D2 VCB would not open. Operations contacted an expert.

Operations prepared D0 and CDF for and access at 9:42 AM.

TeV accesses were allowed to begin at 9:51 AM.

Operations prepared F sector for access at 10:11 AM. This shut off beam to Pbar, BNB, and MTest.

Operations prepared Pbar for access at 10:35 AM. They came out of access an hour later.

A power supply expert reported at 12:37 PM that he had repaired the TeV sector D2 VCB problem.

F sector came out of access at 2:05 PM.

Operations established beam to BNB at 2:42 PM. A Pbar expert began a study period.

At 3:22 PM, experts begin work on TeV RF station #2

Tev RF station #2 repairs were completed at 4:28 PM.

NuMI enclosures were secure at 5:07 PM.

At 5:18 PM the Tev is in a dry squeeze.

At 5:19 PM, problems with power supply LAM61 postponed NuMI target scans until Friday morning.

TeV began a wet squeeze at 7:22 PM.

TeV shot set-up began at 8:05 PM.

At 9:05 PM shot set-up was delayed while experts investigated a problem with injection losses at B0.

Friday July 11

The midnight shift began with the TeV injection loss investigation, with Operations monitoring the stash (380.5E10) and stack (225.8E10), and with BNB taking beam.

Problems with TeV losses at injection were resolved at 12:47 AM.

HEP shot set-up began at 1:30 AM.

Operations established store 6284 at 4:28 AM with an initial luminosity of 286.34E30.

Beam was established to MTest at 5:39 AM.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store. Complete a NuMI target scan and get NUMI back to a standard running mode.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/14/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory